OpenStudy (anonymous):

sinusoidal signal x(n) has a fundamental period of N=10 samples. Determine the smallest angular frequency for which x(n) is periodic.

OpenStudy (anonymous):

eh the fundamental frequency should just be \(f_0=1/T\) where \(T\) is your fundamental period... so \(f_0=1/10\)

OpenStudy (anonymous):

for angular frequency, we're considering a full turn one "cycle" so we should just get \(\omega=2\pi f_0=\pi/5\)

OpenStudy (anonymous):

oh okay cool thanks :) :)